Smooth operations except for high temp alarms due to broken fans.  Replacement fans ordered.   

Site was not getting filled by PanDA for a few weeks, but it's better now. 

Two more NESE gateways added anticipating ramping up.  Working as NESE_DATADISK in AGIS & Rucio.

6PB NESE upgrade arrived, installed, tested, but switches from DELL have been delayed twice. 

Converging on NET2/NESE tape Tier.  Getting helpful feedback from BNL and others in HEP. 

Fred noticed that a few of our oldest nodes were getting a ~50% failure rate, strangely, from stage-out timeouts.  The problem quickly disappeared, but we haven't yet figured out what the cause was. 

User complained about 5 missing files at NET2_DATADISK.  They were indeed missing, marked as gone by DDM.  Don't think it's related to a local issue.